. A sinusoidal wave travels along a string. The time for a particular point to move from


maximum displacement to zero is 0.170 s. What are the (a) period and (b) frequency? (c)


The wavelength is 1.40 m. what is the wave speed?

1
Expert's answer
2022-01-17T09:54:38-0500

"t = 0.17s"

"(a) T = t*4 = 0.17*4 = 0.68s"

"(b) f = \frac{1}{T}= \frac{1}{0.68}= 1.47Hz"

"(c) \lambda = 1.4m"

"\lambda = vT"

"v= \frac{\lambda}{T}= \frac{1.4}{0.68}= 2.06\frac{m}{c}"


"\text{Answer:}"

"(a) T = 0.68s"

"(b) f = 1.47Hz"

"(c) v= 2.06\frac{m}{c}"